OpenStudy (anonymous):

Solve for x: 4 over 5 x + 4 over 3 = 2x

OpenStudy (anonymous):

Fraction within a fraction?

OpenStudy (anonymous):

Or multiplying two fractions?

OpenStudy (anonymous):

Maybe this will help. 4/5 X + 4/3 =2X

OpenStudy (anonymous):

Is this correct?

OpenStudy (anonymous):

Ah, so it's (4/5)(X) + (4/3) = 2X

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Well fractions are messy, so the first thing to do here would be to change the equation so that it no longer has fractions.

OpenStudy (anonymous):

Do you know how to do that?

OpenStudy (anonymous):

No. Fractions is my weakness in algebra

OpenStudy (anonymous):

:) Mine too, you should ask my math teacher. Anyway, a common denominator is needed. Your denominators right now are 5, 3 and 1.

OpenStudy (anonymous):

You can always multiply by 1, and 1 can be anything over itself. You can also multiply individual terms by various versions of "1." Knowing this and that you need a common denominator, what will you do next?

OpenStudy (anonymous):

No. But I do know is that I have to do something to find a common denominator so I can get rid of the fractions.

OpenStudy (anonymous):

I just don't know how to do that step yet.

OpenStudy (anonymous):

Mhm, that something you have to do is multiply by one. :) I'll do one for you if you can tell me what that common denominator can be! HINT: The common denominator is a number that is divisible by all of the denimonators. For example, (1/2) and (1/4) can both be something over 4.

OpenStudy (anonymous):

denominators*

OpenStudy (anonymous):

20 is divisible by the denominators 4 and 5

OpenStudy (anonymous):

Sure, 20 will work. It isn't the lowest common denominator, but it should work.

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

Now how will you change (4X/5) to something over 20?

OpenStudy (anonymous):

Would I multiply?

OpenStudy (anonymous):

Yes! By a version of one. Let me do one for you as an example:

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

Oh!!

OpenStudy (anonymous):

:D Yay! Had an epiphany?

OpenStudy (anonymous):

Haha Yes

OpenStudy (anonymous):

Can you solve it? :)

OpenStudy (anonymous):

Let me check.

OpenStudy (anonymous):

Would it be 4/5?

OpenStudy (anonymous):

Hold on, aren't one of your denominators 3?

OpenStudy (anonymous):

I don't believe 20 is a common denominator if that's true. LOL.

OpenStudy (anonymous):

20 is not nicely divisible by 3. You'd only create another fraction. Unless the question was written wrong.

OpenStudy (anonymous):

Would the common denominator be 15?

OpenStudy (anonymous):

Yes, if the denominators are 5, 3 and 1, the lowest common denominator would be 15. Sorry I didn't catch that sooner.

OpenStudy (anonymous):

It's fine

OpenStudy (anonymous):

So what would the answer be now that the number changed?

OpenStudy (anonymous):

What did you get? I can tell you whether it's right or wrong.

OpenStudy (anonymous):

Sorry Im confused. I forgot what to do when I found the common denominator. I know its 15 but i dont know what to do after that.

OpenStudy (anonymous):

Get all of them to be over 15. Once that's done, the denominators become negligible. I can show you how that works algebraically.

OpenStudy (anonymous):

Ok that would be good.

OpenStudy (anonymous):

Do you know the equation now with the common denominator? Tell me what it is if you do.

OpenStudy (anonymous):

Would it be 4/15 X + 4/15 =2X

OpenStudy (anonymous):

The 2X also needs to be over the same denominator! Only then can we remove the fractions.

OpenStudy (anonymous):

4/15 X + 4/15 = 2X/15

OpenStudy (anonymous):

Doesn't quite work that way. Remember I did one for you? Look at the image again.

OpenStudy (anonymous):

To get that denominator, you have to multiply 4X/5 by 3/3, which is "1." 4X/5 times 3/3. Do you know how to multiply fractions?

OpenStudy (anonymous):

Cross multiply?

OpenStudy (anonymous):

That's one way to do it, but another is to simply multiply across.

OpenStudy (anonymous):

Yeah but now there is a 3 as a denominator

OpenStudy (anonymous):

Yes, and you have to repeat the same process to make that 3 a 15.

OpenStudy (anonymous):

And the 1 that's hiding under the 2X a 15, too.

OpenStudy (anonymous):

I still dont understand what to do.

OpenStudy (anonymous):

Would it be 4/15 X + 4/15 = 2X/15

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

Actually, since you're multiplying both sides by 1/15 it's fine to do that as well, but that won't always work.

OpenStudy (anonymous):

Ah, you aren't though.

OpenStudy (anonymous):

So actually it's completely wrong. You're introducing denominators out of the blue.

OpenStudy (anonymous):

But anyway, I hope that helps. I should do my own work now.

OpenStudy (anonymous):

Do you have the answer?

OpenStudy (anonymous):

You should be able to come up with it yourself.

OpenStudy (anonymous):

Re-read.

OpenStudy (anonymous):

I just dont know what to do.

OpenStudy (anonymous):

I've given you your first step, you should be able to take it from there. Are you opening the images I've sent? Try checking those again. Good luck.